OpenStudy (anonymous):

simplify the square root of 63a2b2c

Nnesha (nnesha):

hint :\[\huge\rm \sqrt{x} = x^{\frac{ 1 }{ 2 }}\]

OpenStudy (anonymous):

Another hint: sqrt(63) = 3*sqrt(7)

Nnesha (nnesha):

this hint is better than the previous one ;P\[\huge\rm \sqrt[n]{x^m} = x^{\frac{ m }{ n }}\]